40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7247     Entry time: Wed Aug 22 01:54:03 2012
Author: Jenne 
Type: Update 
Category: General 
Subject: List o' things: YarmASS, ETMXTcamera, POXwhiteningTriggering 

While meditating on other things, I have noticed / found the following today:

YARM ASS works okay.  Yesterday I measured the sensing matrix for the ASS for both arms (although I forgot to copy one of the matrix elements to my text file for Xarm - needs remeasuring).  I put the Yarm matrix in (after appropriate inversion, only non-zero pitch-to-pitch, yaw-to-yaw elements).  I turned on the Yarm ASS, and  the yaw converged pretty quickly (couple of minutes), with gains of -1 in the servos, overall gain of anywhere between 0.005 and 0.010.  The pitch took much longer, and I had to 'pause' several times by turning off the overall gain for the yarm ass when the MC lost lock (which has happened several times tonight - unknown cause).  Eventually, the pitch settled out, and quit changing, but the lockin outputs weren't zero, even though the error signal for the servos were almost zero (gains for the pitch servos were -0.5, overall gain ~0.005 was better than 0.01 - higher gain caused oscillations in the lockin outputs).  I think this means that I need to remeasure the yarm pitch ass matrix.  It's still much, much faster to just turn on the dithers, watch the striptool of the lockin outputs, and align the cavity by hand.

I think the ETMX Trans camera view is clipped a little bit.  I went down there, and it doesn't seem to be on the last optic before the camera, and moving the spot on the camera doesn't change the shape of the image, so I don't think it's on the camera.  We should look into this, since it's either clipping on the BS that separates some camera beam from the TRX beam, or TRX is getting a clipped beam too.  If the clipping is any earlier in the Trans path, the Trans QPD could also have some clipping.  This requires investigation.  The xarm trigger needs to be reset/disabled so we don't lose lock every time we block the TRX beam (as was happening to me).

XARM really doesn't like to relock unless the POX whitening is turned off.  Good flashes, doesn't really catch (10+ min waiting (while working on Yarm stuff) ).  After turning off the whitening, it catches almost immediately. Even though it's on the to-do list to rethink the tuning of our whitening, we should probably implement the whitening triggering now anyway.  It'll make things easier.

The double integrator that Rana implemented in the X and Y arm servo filters last week take 8 seconds to turn off (due to Foton settings), so even though they are triggered to turn off immediately upon lockloss, they sit around and integrate for 8 seconds, so have huge signals.  If the cavity flashes and the locking trigger engages during that 8 seconds, we send a huge kick to the ETMs.  I'm modeling the response of the filters to an impulse and noise, particularly in the case of ramping on the double integrators.  The problem is that a flat filter has 0deg phase, but the double integrator has 180deg phase at low frequencies, so there's some weird sign flipping that can happen as we ramp - this is part of what I'm modeling.

MC is losing lock unusually often tonight.  Everything on the servo board screen looks normal (which is good since that's all set by the autolocker).  I just disabled the test exc in, but that's been left enabled for a while now, and it hasn't (I think?) been a problem since there shouldn't be anything connected to the board there.  PMC transmission is a little low, 0.816, and FSS is starting to get near -1 on the slow actuator adjust, but we've seen locking of the PMC problems around -1.5 or -2 of the FSS, and the adjust value was at -0.8 earlier tonight and we still had MC locking problems.  I have had the seismic channels open on Dataviewer for the last several hours, and I'm not seeing any spikes in any of the Guralp channels which correspond to the times that the MC loses lock.  BLRMS don't seem particularly high, so MC lockloss cause is still a mystery for today.

The ETMX monitor selector on the VIDEO screen seems not to be switching the actual camera that's shown on the monitor.  Using the script command itself works, so my screen is wrong.

ELOG V3.1.3-